### v1

@@ 3.1 Preliminaries
    Abduction: Infer $P(\boldsymbol{\epsilon}|\mathbf{x})$, the state of the world (exogenous noise) that is compatible with the observation $\mathbf{x}$.
@@
    Action: Replace the structural equations $do(x_{i}=y)$ corresponding to the intervention, resulting in a modified SCM $\widetilde{\mathcal{G}}:=\mathcal{G}_{\mathbf{x};do(x_{i}=y)}=(\widetilde{\mathbf{S}},P(\boldsymbol{\epsilon}|\mathbf{x})$.
@@
    Prediction: Use the modified model to compute $P_{\widetilde{\mathcal{G}}}(\mathbf{x})$.
Mechanisms In the context of counterfactual image generation, we examine the effect of a change to a parent variable (attribute) on the image. From now on, we will use the notation $x$ and $x^{*}$ for the factual and counterfactual image respectively, and $pa_{x}$, $pa^{*}_{x}$ for the factual and counterfactual attributes.
The conventional linear mechanisms that have been used in the statistical literature for deriving SCMs are not useful when modelling high-dimensional variables, such as images [40]. To overcome this limitation, end-to-end deep learning techniques have been incorporated in the field of causal inference [8]. The framework we are examining, called Deep-SCM [31] is a prime example of such innovation. Introducing invertible mechanisms for all causal variables, enables tractable counterfactual inference by following the aforementioned Abduction-Action-Prediction paradigm.
Three categories of invertible mechanisms are introduced:
  1. 1.
    Invertible, explicit used for the attribute mechanisms in conjunction with Conditional Normalising Flows [46]. This mechanism is invertible by design.
  2. 2.
    Amortised, explicit employed for high-dimensional variables, such as images. This can be developed with Conditional VAE s [20, 16] or extensions like Conditional Hierarchical Variational Autoencoder (HVAE) [21, 41]. The structural assignment is decomposed into a low-level invertible component, in practice the reparametrisation trick, and a high-level non-invertible component that is trained in an amortized variational inference manner, which in practice can be a probabilistic convolutional decoder. This also entails a noise decomposition of $\epsilon_{i}$ into two noise variables $(\epsilon,z)$, s.t. $P(\epsilon_{i})=P(\epsilon)P(z)$.
  3. 3.
    Amortised, implicit This mechanism can also be employed for the imaging variables. However, it does not rely on approximate maximum-likelihood estimation for training, but optimises an adversarial objective with a conditional implicit-likelihood model. It was introduced in [3] and developed further with Conditional GANs [28, 7].
3.2 Methods Considered in the Benchmark
Conditional Normalising Flows To enable tractable abduction of $pa_{x}$, invertible mechanisms can be learned using conditional normalising flows. A normalising flow maps between probability densities defined by a differentiable, monotonic, bijection. These mappings are a series of relatively simple invertible transformations and can be used to model complex distributions [44]. The mechanism for each attribute ($pa_{x}$) is a normalising flow $f(\epsilon)$. The base distribution for the exogenous noise is typically assumed to be Gaussian $\epsilon\sim N(0,I)$. The distribution of $pa_{x}$ can be computed as $P(pa_{x})=P(\epsilon)\left|\text{det}\,\nabla f(\epsilon)\right|^{-1}$. Thus, the conditional distribution $P(pa_{x}|pa_{pa_{x}})$ can be computed as:
@@ 3.2 Methods Considered in the Benchmark
Conditional VAE To produce counterfactuals via leveraging conditional VAE s [20, 16], the high-dimensional variable $x$ along with its factual parents $pa_{x}$ are encoded into a latent representation $z\sim q_{\phi}(z|x,pa_{x})$ which is sampled through the reparameterisation trick: $z=\mu_{\phi}(x,pa_{x})+\sigma_{\phi}(x,pa_{x})*\epsilon$, where $\epsilon\sim N(0,I)$ [4]. Then, the latent variable $z$ and the counterfactual parents $pa^{*}_{x}$ are decoded to obtain the counterfactual distribution $(\mu_{\theta}^{*},\sigma_{\theta}^{*})$. Since the observational distribution is assumed to be Gaussian, sampling can be performed by forwarding the decoder with the factual parents and then applying a reparametrization: $x=\mu_{\theta}(z,pa_{x})+\sigma_{\theta}(z,pa_{x})*\epsilon$, $\epsilon\sim N(0,I)$. Thus, the exogenous noise can be expressed as $\epsilon=\frac{x-\mu_{\theta}(z,pa_{x})}{\sigma_{\theta}(z,pa_{x})}$. Finally, this $\epsilon$ is used to sample from the counterfactual distribution as follows: $x^{*}=\mu_{\theta}(z,pa^{*}_{x})+\sigma_{\theta}(z,pa^{*}_{x})*\epsilon$.
Conditional HVAE A conditional structure of the top-down HVAE [41, 2, 47] can be adopted to improve the quality of produced samples and abduct noise more accurately. As we consider Markovian SCMs, the exogenous noise $\boldsymbol{\epsilon}$ has to be independent, therefore as the latents $z_{1:L}$ are downstream components of the exogenous noise, the corresponding prior $p_{\theta}(z_{1:L})$ has to be independent from $pa_{x}$. In order to accomplish this, the authors of [4] propose to incorporate $z_{i}$ and $pa_{x}$ at each top-down layer through a function $f^{i}_{\omega}$ which is modelled as a learnable projection network. Particularly, the output of each top-down layer can be written as:
@@
Conditional GAN Finally, we also consider the conditional GAN-based framework [3] for counterfactual inference. This setup includes an encoder $E$, that given an image $x$ and its parent variables $pa_{x}$ produces a latent code $z_{x}=E(x,pa_{x})$. The generator $G$ then receives either a noise latent $z$ or the produced $z_{x}$ together with the parents $pa_{x}$ to produce an image $x^{\prime}=G(z^{\prime},pa_{x})$. The image alongside the latent and the parents $pa_{x}$ are then fed into the discriminator. The encoder and the generator are trained together as in [7] to deceive the discriminator in classifying generated samples from a noise latent $z$ as real, while classifying samples produced from $z_{x}$ (real image latents) as fake, while the discriminator has the opposite objective. Providing $pa_{x}$ as input to all models is proven to make conditioning on parents stronger. The conditional GAN is optimised as follows:
@@
$x^{*}=G(E(x,(pa_{x})),pa^{*}_{x}),$ (6)
However, as it is discussed in the original paper [3] and as we confirmed empirically, this objective is not enough to enforce accurate abduction of the noise, resulting in low reconstruction quality. In order to alleviate this, the encoder is fine-tuned, while keeping the rest of the network fixed, to minimise explicitly the reconstruction loss on image and latent space. This cyclic cost minimisation approach, introduced by [6], enables accurate learning of the inverse function of the generator.
3.3 Evaluation Metrics
The evaluation of counterfactual inference has been formalised through the axioms of [10, 12]. The authors of [30] utilise such an axiomatic definition to introduce three metrics for evaluating image counterfactuals: Composition, Effectiveness, and Reversibility  [^3] We did not include reversibility, since the applied interventions cannot always be cycle-consistent. This was also the case in the follow up work by the same authors [4].. While it is necessary for counterfactual images to respect these axiom-based metrics, we find that they are not sufficient for a perceptual definition of successful counterfactuals. The additional desiderata we consider are realism of produced images, as well the minimality or sparseness of changes. For the former, we employ the FID metric, whilst for the latter, we employ the Counterfactual Latent Divergence metric proposed in [37]. All adopted metrics do not require access to ground truth counterfactuals.
Composition If we force a variable $X$ to a value $x$ it would have without the intervention, it should have no effect on the other variables [10]. It follows then, that we can have a null-intervention that should leave all variables unchanged. Based on this property one can apply the null-intervention $m$ times, denoted as $f^{m}_{\emptyset}$, and measure the $l_{1}$ distance $d(\cdot)$ between the $m$ produced and the original image [30]. To produce a counterfactual under the null-intervention, we abduct the exogenous noise and skip the action step.
@@ 3.3 Evaluation Metrics
As the $\ell_{1}$ distance on image space yields results that do not align with our perception and favours shortcut features such as simpler backgrounds, we extend Composition by computing the $\ell_{1}$ norm of VGG-16 embeddings and the Learned Perceptual Similarity (LPIPS) [53].
Effectiveness aims to identify how successful is the performed intervention. In other words, if we force a variable $X$ to have the value $x$, then $X$ will take on the value $x$ [10]. In order to quantitatively evaluate effectiveness for a given counterfactual image we leverage an anti-causal predictor $g^{i}_{\theta}$ trained on the data distribution, for each parent variable $pa^{i}_{x}$  [30]. Each predictor, then, approximates the counterfactual parent $pa^{i*}_{x}$ given the counterfactual image $x^{*}$ as input
@@
where $d(\cdot)$ is the corresponding distance, defined as a classification metric for categorical variables and as a regression metric for continuous ones.
Realism For realism we use the Fréchet Inception Distance (FID) [15] which captures the similarity of the set of counterfactual images to all images in the dataset. Real and counterfactual samples are fed into an Inception v3 model [43] trained on ImageNet, in order to extract their feature representations, that capture high-level semantic information.
Minimality The need for minimality of counterfactuals can be justified on the basis of the sparse mechanism shift hypothesis [39]. While we can achieve an approximate estimate of minimality by combining the composition and effectiveness metrics, we find that an additional metric is helpful to measure the closeness of the counterfactual image to the factual. For this reason, we leverage a modified version of the Counterfactual Latent Divergence (CLD) metric introduced in [37]:

### v5

@@ A.1 Mechanisms and Models
    Invertible, explicit used for the attribute mechanisms in conjunction with Conditional Normalising Flows (53). This mechanism is invertible by design.
@@
    Amortised, explicit employed for high-dimensional variables, such as images. This can be developed with Conditional VAE s (25; 20) or extensions like Conditional Hierarchical Variational Autoencoder (HVAE) (26; 47). The structural assignment is decomposed into a low-level invertible component, in practice the reparametrisation trick, and a high-level non-invertible component that is trained in an amortized variational inference manner, which in practice can be a probabilistic convolutional decoder. This also entails a noise decomposition of $\epsilon_{i}$ into two noise variables $(\epsilon,z)$, s.t. $P(\epsilon_{i})=P(\epsilon)P(z)$.
@@
    Amortised, implicit This mechanism can also be employed for the imaging variables. However, it does not rely on approximate maximum-likelihood estimation for training, but optimises an adversarial objective with a conditional implicit-likelihood model. It was introduced in (6) and developed further with Conditional GANs (33; 10).
The models we use to implement these mechanisms are detailed below.
Conditional Normalising Flows To enable tractable abduction of $pa_{x}$, invertible mechanisms can be learned using conditional normalising flows. A normalising flow maps between probability densities defined by a differentiable, monotonic, bijection. These mappings are a series of relatively simple invertible transformations and can be used to model complex distributions (51).
In particular, a random variable $z_{0}$ is mapped through a normalising flow $f$, to a transformed variable $z_{K}=f_{K}(f_{K-1}(\ldots f_{1}(z_{0})))$ and its distribution $p(z_{K})=p(z_{0})-\sum_{k=1}^{K}\left|\text{det}\,df_{k}(z_{k-1})\right|\,dz_{k-1}$, where $p(z_{0})$ is a standard normal base distribution and each $f_{k}$ is a simple, monotonically increasing function which has a closed form with an easy to calculate derivative.
The mechanism for each attribute ($pa_{x}$) is a normalising flow $f(\epsilon)$. The base distribution for the exogenous noise is typically assumed to be Gaussian $\epsilon\sim N(0,I)$. The distribution of $pa_{x}$ can be computed as $P(pa_{x})=P(\epsilon)\left|\text{det}\,\nabla f(\epsilon)\right|^{-1}$. Thus, the conditional distribution $P(pa_{x}|pa_{pa_{x}})$ can be computed as:
@@
Conditional VAE Conditional VAE s (25; 20) are composed of a generative model $p_{\theta}(x,z|pa_{x})=p_{\theta}(x|z,pa_{x})p(z)$ and an inference model $q_{\phi}(z|x,pa_{x})$, where $pa_{x}$ is the conditioning signal. The likelihood $p_{\theta}(x|z,pa_{x})$ and the approximate posterior $q_{\phi}(z|x,pa_{x})$ are assumed to be diagonal Gaussians, while the prior $p(z)$ is fixed as standard Gaussian $p(z)\sim N(0,I)$. The evidence lower bound loss can be expressed as:
$\text{ELBO}_{\beta}(\phi,\theta)=\mathbb{E}_{z\sim q_{\phi}(z|x,pa_{x})}[p_{\theta}(x|z,pa_{x})]-\beta D_{KL}[q_{\phi}(z|x,pa_{x})||p(z)]$ (2)
The distributions $q_{\phi}(z|x,pa_{x})$, $p_{\theta}(x|z,pa_{x})$ are parameterised as neural networks with parameters $\phi$ and $\theta$ accordingly that output the $\mu$ and $\sigma$ of the distributions. The ELBO is jointly maximised with respect to $\phi$, $\theta$.
To produce counterfactuals via leveraging conditional VAE s (25; 20), the high-dimensional variable $x$ along with its factual parents $pa_{x}$ are encoded into a latent representation $z\sim q_{\phi}(z|x,pa_{x})$ which is sampled through the reparameterisation trick: $z=\mu_{\phi}(x,pa_{x})+\sigma_{\phi}(x,pa_{x})*\epsilon$, where $\epsilon\sim N(0,I)$ (7). Then, the latent variable $z$ and the counterfactual parents $pa^{*}_{x}$ are decoded to obtain the counterfactual distribution $(\mu_{\theta}^{*},\sigma_{\theta}^{*})$. Since the observational distribution is assumed to be Gaussian, sampling can be performed by forwarding the decoder with the factual parents and then applying a reparametrization: $x=\mu_{\theta}(z,pa_{x})+\sigma_{\theta}(z,pa_{x})*\epsilon$, $\epsilon\sim N(0,I)$. Thus, the exogenous noise can be expressed as $\epsilon=\frac{x-\mu_{\theta}(z,pa_{x})}{\sigma_{\theta}(z,pa_{x})}$. Finally, this $\epsilon$ is used to sample from the counterfactual distribution as follows: $x^{*}=\mu_{\theta}(z,pa^{*}_{x})+\sigma_{\theta}(z,pa^{*}_{x})*\epsilon$.
Conditional HVAE Hierarchical VAE s (26; 47) extend classical VAE s by introducing $L$ groups of latent variables $\boldsymbol{z}=\{z_{1},z_{2},...,z_{L}\}$. One widely used variation of the HVAE is the top-down VAE which is proposed in (47). In the top-down approach (47; 5; 55) both the prior and the approximate posterior emit the groups of the latent variables in the same top-down order, thus the two distributions can be factorised as:
$p_{\theta}(z_{1:L})=p_{\theta}(z_{L})p_{\theta}(z_{L-1}|z_{L})...p_{\theta}(z_{1}|z_{i>1}),$ (3)
$q_{\phi}(z_{1:L}|x)=q_{\phi}(z_{L}|x)q_{\phi}(z_{L-1}|z_{L},x)...q_{\phi}(z_{1}|z_{i>1},x)$ (4)
A conditional structure of the top-down HVAE (47; 5; 55) can be adopted to improve the quality of produced samples and abduct noise more accurately. As we consider Markovian SCMs, the exogenous noise $\boldsymbol{\epsilon}$ has to be independent, therefore as the latents $z_{1:L}$ are downstream components of the exogenous noise, the corresponding prior $p_{\theta}(z_{1:L})$ has to be independent from $pa_{x}$. In order to accomplish this, the authors of (7) propose to incorporate $z_{i}$ and $pa_{x}$ at each top-down layer through a function $f^{i}_{\omega}$ which is modelled as a learnable projection network. Particularly, the output of each top-down layer can be written as:
@@
Counterfactual training This fine-tuning step was proposed in (7) to alleviate the ignoring of the conditioning that happened for some datasets. To achieve this, the model generates counterfactuals and utilises the losses of classifiers trained on observational data, while keeping the reconstruction loss above a threshold. In brief, the lowest ELBO loss is incorporated early during training into the counterfactual loss by introducing a Lagrange coefficient, thus transforming the setup into a Lagrangian optimisation problem (41). The overall objective can be expressed as (7):
$L(\theta,\phi,\lambda;\mathbf{x},\mathbf{pa_{x}})=-\sum_{k=1}^{K}\mathbb{E}_{\begin{subarray}{c}\mathbf{pa^{*}_{x_{k}}}\sim p(\mathbf{pa_{x_{k}}})\\ \mathbf{x^{*}}\sim P_{\mathbf{CF}}(\mathbf{x^{*}}\mid\text{do}(\mathbf{pa^{*}_{x_{k}}}),\mathbf{x})\end{subarray}}\left[\log q_{\psi_{k}}(\mathbf{pa^{*}_{x_{k}}}\mid\mathbf{x^{*}})\right]$
$-\lambda(c-ELBO(\theta,\phi;\mathbf{x},\mathbf{pa_{x}})),$ (8)
where $q_{\psi_{k}}$ is the $k_{th}$ parent classifier, $c$ is the ELBO that the model achieves during pretraining and $P_{CF}$ the counterfactual distribution. To optimize the above objective gradient descent is performed on the HVAE parameters $\phi,\theta$ and gradient ascent on the Lagrange coefficient $\lambda$.
It should be noted that this process can introduce bias from the classifiers, which bias is in any case present in the data and can be picked up from the model even without this step. Finally, using the same classifiers to measure effectiveness is unfair compared to the other models, which we alleviate by using different classifiers to guide the counterfactual finetuning.
Conditional GAN Finally, we also consider the conditional GAN-based framework (6) for counterfactual inference. This setup includes an encoder $E$, that given an image $x$ and its parent variables $pa_{x}$ produces a latent code $z_{x}=E(x,pa_{x})$. The generator $G$ then receives either a noise latent $z$ or the produced $z_{x}$ together with the parents $pa_{x}$ to produce an image $x^{\prime}=G(z^{\prime},pa_{x})$. The image alongside the latent and the parents $pa_{x}$ are then fed into the discriminator. The encoder and the generator are trained together as in (10) to deceive the discriminator in classifying generated samples from a noise latent $z$ as real, while classifying samples produced from $z_{x}$ (real image latents) as fake, while the discriminator has the opposite objective. Providing $pa_{x}$ as input to all models is proven to make conditioning on parents stronger. The conditional GAN is optimised as follows:
@@
$x^{*}=G(E(x,(pa_{x})),pa^{*}_{x})),$ (10)
However, as it is discussed in the original paper (6) and as we confirmed empirically, this objective is not enough to enforce accurate abduction of the noise, resulting in low reconstruction quality. In order to alleviate this, the encoder is fine-tuned, while keeping the rest of the network fixed, to minimise explicitly the reconstruction loss on image and latent space. This cyclic cost minimisation approach, introduced by (9), enables accurate learning of the inverse function of the generator.
Error in the image space $\begin{array}[]{l}\text{Error in the image space }L_{x}=\mathbb{E}_{x\sim q(x)}\lVert x-G(E(x,pa_{x}),pa_{x})\rVert_{2}\\ \text{Error in the latent space }L_{z}=\mathbb{E}_{z\sim p(z)}\lVert z-E(G(z,pa_{x}),pa_{x})\rVert\end{array}$ (11)
A.2 Metrics
Composition If we force a variable $X$ to a value $x$ it would have without the intervention, it should have no effect on the other variables (13). It follows then, that we can have a null-intervention that should leave all variables unchanged. Based on this property one can apply the null-intervention $m$ times, denoted as $f^{m}_{\emptyset}$, and measure the $l_{1}$ distance $d(\cdot)$ between the $m$ produced and the original image (35). To produce a counterfactual under the null-intervention, we abduct the exogenous noise and skip the action step.
@@ A.2 Metrics
As the $\ell_{1}$ distance on image space yields results that do not align with our perception and favours shortcut features such as simpler backgrounds, we extend composition by computing the Learned Perceptual Similarity (LPIPS) (63).
Effectiveness aims to identify how successful is the performed intervention. In other words, if we force a variable $X$ to have the value $x$, then $X$ will take on the value $x$ (13). In order to quantitatively evaluate effectiveness for a given counterfactual image we leverage an anti-causal predictor $g^{i}_{\theta}$ trained on the data distribution, for each parent variable $pa^{i}_{x}$  (35). Each predictor, then, approximates the counterfactual parent $pa^{i*}_{x}$ given the counterfactual image $x^{*}$ as input
@@
where $d(\cdot)$ is a distance, defined as a classification metric for categorical variables and as a regression metric for continuous ones.
Realism For realism we use the Fréchet Inception Distance (FID) (19) which captures the similarity of the set of counterfactual images to all images in the dataset. Real and counterfactual samples are fed into an Inception v3 model (50) trained on ImageNet, in order to extract their feature representations, that capture high-level semantic information.
We denote as $q$ the distribution of the feature representations of counterfactuals and as $p$ the distribution of the feature representations of the factuals. Fréchet distance $d$ is defined as the distance between the Gaussian with mean $(m_{q},C_{q})$ obtained from $q$ and the Gaussian with mean $(m_{p},C_{p})$ obtained from $p$. FID is defined as:
$d^{2}((m_{q},C_{q}),(m_{p},C_{p}))=||m_{q}-m_{p}||_{2}^{2}+\text{Tr}(C_{q}+C_{p}-2(C_{q}C_{p})^{\frac{1}{2}}).$ (14)
Minimality The need for minimality of counterfactuals can be justified on the basis of the sparse mechanism shift hypothesis (46). While we can achieve an approximate estimate of minimality by combining the composition and effectiveness metrics, we find that an additional metric is helpful to measure the closeness of the counterfactual image to the factual. For this reason, we leverage a modified version of the Counterfactual Latent Divergence (CLD) metric introduced in (44):